# MobCLIP: Learning General-purpose Geospatial Representation at Scale

### v3

@@ Abstract
Representation learning of geospatial locations remains a core challenge in achieving general geospatial intelligence. Current embedding methods often lack versatility, limiting their utility across diverse tasks in both human and natural domains. We present MobCLIP, the first nationwide general-purpose location encoder, integrating an unprecedented diversity of data modalities through effective and scalable multimodal fusion. Adopting a novel CLIP-based architecture, our framework aligns 100M+ POIs, nationwide remote sensing imagery, and structured demographic statistics with a billion-edge mobility graph. By tokenizing spatial locations into grid cells inspired by Vision Transformers, we establish a unified representation space bridging mobility patterns and multimodal features. To rigorously evaluate the general-purpose effectiveness of MobCLIP, we construct a benchmark dataset composed of 11 downstream prediction tasks across social, economic, and natural domains. Experiments show that MobCLIP, with four input modalities and a compact 128-dimensional representation space, achieves significantly superior general-purpose predictive performances than state-of-the-art models by an average of 35%. Thanks to the effective integration of human-centric modalities, the performance gain is particularly profound in human-centric tasks, such as energy consumption (+260%), offline retail consumption amount (+98%), and crime cases (+95%) predictions. Echoing LLM scaling laws, we further demonstrate the scaling behavior in geospatial representation learning. We open-source code and pretrained models at: https://github.com/ylzhouchris/MobCLIP.
@@ 1 Introduction
Geospatial representation learning is crucial for enabling AI systems to understand and reason about spatial patterns, relationships, and dynamics across diverse geographic contexts [6, 54, 28]. This process involves encoding regions into low-dimensional vectors that capture both their intrinsic characteristics and their correlations. Despite some early attempts [43, 20, 39], the lack of large-scale general-purpose region representations remains a fundamental bottleneck.
There has been a rise in geospatial representation learning studies, attempting to construct pretrained information-dense spatial representations for specific downstream tasks. Wu et al. [42] established a unified location encoding framework that consolidates 15 commonly recognized location encoders that mainly come in two folds. Position encoding functions [26] maps 2D coordinates into high-dimentional feature space using ideas such as spherical harmonics [26] and sinusoidal activations [33]. Early attempts of self-supervised location embeddings methods primarily combine image data, either geo-tagged photos or satellite imagery, with geospatial locations using contrastive learning models [38, 25, 19]. These methods demonstrate the potential of "clipping" contextual information to geospatial regions, achieving success in downstream tasks like global localization [38].
However, geospatial inference tasks are inherently multimodal in nature. Effective and scalable representation learning requires careful selection, alignment, and integration of multiple informative input modalities. Mobility data, for instance, captures human behaviors in cities, with mobility graphs enhancing correlations between locations through movement patterns. While studies have shown the effectiveness of mobility graphs at city scales [56, 46, 51, 52], demographic statistics have also proven valuable in urban representation learning [40]. Among these attempts, the contrastive learning framework and the CLIP [31] model has become prominent in multimodality alignment. The difficulty of effective multimodal fusion increases dramatically with the number of input modalities. Multimodal alignment for geospatial data remains a challenge.
Geospatial representations, if learned well, should demonstrate superior general-purpose inference capabilities in a variety of downstream tasks spanning social, economic, and natural prediction domains. Existing methods [26, 19] focus primarily on natural tasks such as forest coverage estimation, elevation estimation, and bird species classifications [25], paying insufficient attention to human-centric tasks related to the social-economic dynamics essential in global cities. There is a lack of comprehensive benchmarks involving both human-centric and natural tasks, making it difficult to evaluate the general-purpose inference capabilities of location encoders.
To address these challenges, we propose MobCLIP, the first nationwide general-purpose location encoder that integrates an unprecedented diversity of data modalities through a scalable multimodal fusion framework. Unlike traditional methods, MobCLIP innovatively leverages mobility data as the backbone, fusing information from other modalities into it using a CLIP-based approach. This strategy not only improves multimodal integration but also explicitly addresses the traditional underemphasis on human-centric data. Our contributions can be summarized as follows:
  • We introduce MobCLIP, the first nationwide, general-purpose pretrained location encoder, designed for diverse tasks beyond traditional geospatial domains.
  • We introduce the novel use of mobility data as the backbone, combined with a self-supervised multimodal pretraining framework, thereby emphasizing the significance of human-centric information in geospatial learning.
@@
  • We construct a benchmark dataset of 11 downstream prediction tasks spanning social, economic, and natural domains for performance evaluation. Extensive experiments show that our model achieves significantly superior performance—improving by over 35% on average—across all general-purpose tasks. Performance gains are especially profound in human-centric tasks.
  • We open-source our methodology, a distilled version of MobCLIP, and a benchmark dataset for general-purpose evaluation to advance research on general-purpose pretrained location encoders.
@@ 2 Related Works
Pretrained foundation model for general geospatial inference.
Geospatial data is inherently diverse and multimodal, comprising various forms such as maps [41, 2, 8], texts [21, 22, 11], images [10, 14], and graphs [43, 39], all of which have been widely employed to advance the understanding of urban areas. To effectively leverage this diversity, multimodal region representation learning has emerged as a paradigm for integrating cross-modality information in a self-supervised manner to learn task-agnostic regional embeddings [47, 53, 52, 51]. However, such frameworks require each city to be trained individually while heavily relying on the availability of multimodal data in the study area, resulting in limited generalizability. Following the success of pretrained Foundation Models (FMs) in language and vision domains, there is increasing interest and active discussion around developing foundation models for geospatial artificial intelligence (GeoAI) [27, 54]. ReFound [45], one of the early attempts, aims to distill extensive knowledge from general FMs while incorporating domain-specific objectives for enhanced performance. It collects POI and satellite image data from five major cities in China as the pretraining corpora, which is relatively limited in scope and diversity of data modalities compared to ours. Perhaps the one most similar to us in concept is the Population Dynamics Foundation Model (PDFM) [1], which encodes postal codes and counties across the United States into embeddings by aggregating features such as busyness and search trends via a neighborhood-based graph. However, comparatively, it still falls short in the richness and diversity of modalities. Additionally, the model adopts simple concatenation for cross-modality fusion and relies solely on geographical proximity for information aggregation, missing the higher-level correlations captured by our mobility graph.
@@
A straightforward approach for representing geolocations involves using analytical transformations to convert 2D or 3D coordinate inputs into higher-dimensional vectors while ensuring that the resulting embeddings preserve the geographical distances between locations. For instance, Space2Vec [24] encodes locations using sinusoid functions with different frequencies in a 2D Euclidean space. Extending from 2D space, Neural Radiance Fields (NeRF) [29] encodes continuous 3D coordinates via Fourier input mapping. To further account for the Earth’s spherical geometry, Sphere2Vec [26] and Siren [33] introduce specialized position encoders designed to capture spherical distances between coordinates. The position encoders are typically followed directly by a neural network, which is trained in a supervised manner on task-specific datasets.
Self-supervised location representation pretraining.
To reduce reliance on heavily labeled data, which is both scarce and costly to obtain, and to leverage the abundance of unlabeled geo-tagged and satellite images [49, 50, 48], self-supervised pretraining methods have emerged for location representation learning. Contrastive Spatial pretraining (CSP) [25] learns location embeddings from unlabeled imagery using a contrastive objective. CSP is pretrained, fine-tuned, and tested on the same datasets (i.e., iNat2018 [37] and fMoW [9]), without demonstrating generalizability to other tasks. SatCLIP [19] and GeoCLIP [38] both pretrain location encoders via contrastive learning on coordinate-image pairs to generate global location embeddings. Their primary differences lie in the architectures of their location and image encoders, as well as the pretraining datasets used. Although existing location-based pretraining methods excel in enabling global-scale embedding queries, they fail to incorporate other geospatial modalities beyond images and overlook crucial factors, such as human activities, that are essential for learning general-purpose location representations.
@@ 3 Methodology
The framework is designed to learn versatile region representations by integrating geospatial features from four distinct modalities: human mobility (modeled as a graph), Points-of-Interest (POIs) as textual data, satellite imagery as visual input, and demographic distributions represented as tabular categorical histograms. Figure 1 illustrates the proposed methodological framework.
Figure 1: The methodological framework of MobCLIP.
@@ 3.1 Mobility as the Backbone
Geographic location serves as a precise anchor for aligning multimodal data. Prior region embedding methods [42, 19] often rely on coordinates to fuse different modalities, but this limits the inclusion of network-based signals. Human activity frequently spans geographic boundaries via transport or digital networks, and mobility data—represented as links in a spatial graph—captures these non-local patterns. We propose to use the mobility graph as a backbone for multimodal fusion in geospatial representation learning.
@@
Raw spatial interactions are sparse and misaligned in time and space, making coordinate-based graphs too fragmented to reflect major flows. Inspired by Vision Transformers (ViT) [18], which partition images into patches, we grid geographic space and build graphs at the cell level. We adopt the H3 grid [36], which minimizes distortion compared to Geohash and Google S2 [3], offering balanced global coverage. At level 6 ($\sim$36 km² per cell), we construct a nationwide mobility graph from real-world transaction data as presented in the “Graph Construction” section in Figure 1 . This large-scale network captures human movement patterns and serves as the backbone for multimodal alignment and representation learning.
@@
We utilize LINE (Large-scale Information Network Embedding) [34] to preprocess the full mobility graph and generate 128-dimensional embeddings for each H3 grid cell. Based on empirical evaluation, we select LINE embeddings with second-order proximity, which offer superior performance. When paired with the graph sampling strategy described below, this approach effectively preserves the information from the full mobility graph with minimal computational effort.
@@
Graph Neural Networks (GNNs) have proven to be effective frameworks for capturing relationships between nodes, taking into account high-order dependencies. GNNs operate through a message-passing mechanism, where each node updates its representation by aggregating information from its neighboring nodes. To encode a graph $\mathcal{G}$ using $L$ layers of GNNs, the $l^{th}$ layer embedding of a node $i$ is computed as follows:
@@
where $\mathcal{N}_{i}$ refers to the neighborhood of the node $i$, and $\textit{AGG}(\cdot)$ is the aggregation function, which can vary across different models. In our framework, we employ LightGCN [12] as our mobility graph encoder. It simplifies the traditional GNN architecture by removing unnecessary feature transformation and nonlinear activation, and only keeping the most essential component: neighborhood aggregation. In LightGCN, node features are initialized as learnable parameters and by utilizing $L$ layers of propagation, it generates $L+1$ distinct embeddings for each node, namely $(e^{(0)},e^{(1)},\dots,e^{(L)})$:
@@
We use the pre-encoded LINE embeddings to initialize the node features and construct a mobility subgraph as the graph encoder input. This subgraph is extracted from the full graph using a top-$k$ sampling strategy based on a percentage threshold, ensuring that the links with the highest flows are retained for each node. This strategy aims to address the "long-tail" distribution patterns commonly seen in geospatial studies, where a small number of regions experience very high flows, while the majority of regions exhibit significantly lower flows that have minimal impact on the overall graph structure. We adopt a top sampling ratio of 10% for the main experiments.
@@ 3.2 Auxiliary Multimodality Encoder
For general-purpose geospatial inference, the selection of fundamental input modalities that are (1) effective for a variety downstream tasks in both human-centric and nature domains, and (2) complementary to each other, is critical for the performance multimodal representation learning. Based on existing knowledge from spatial data science and urban studies and demonstrated effectiveness in literature, we select three fundamental auxiliary modalities to complement the mobility backbone, including Point-of-Interest (POI), Demographic Statistics, and Satellite imagery.
Different unimodal encoders are selected to encode auxiliary modalities. (1) Point-of-Interest (POI): We encode both POI categorical statistics and textual information using language models [7]. Specifically, we use the BGE-m3 model [5] to encode both POI names and categories. Embedding of all POIs located within each H3 grid was then averaged to be the grid embedding. (2) Satellite imagery: We use an off-the-shelf pretrained image encoder, RemoteClip [23], to generate image embeddings for each H3 grid. (3) Demographic Statistics: The categorical demographic data in tabular form is encoded using an MLP-based encoder.
@@ 3.3 Multimodality Alignment with Mobility
We adopt a CLIP-inspired contrastive learning scheme to jointly project embeddings of the four modalities into a joint embedding space and aligned under a contrastive learning objective.
We train our model using a CLIP-inspired contrastive learning scheme, in which mobility serves as the central modality that other modalities align to. We construct modality tuples $(\mathcal{M},\mathcal{I},\mathcal{T},\mathcal{D})$ for all grids in the study area, where $\mathcal{M}$ indicates mobility, and $\mathcal{I}$, $\mathcal{T}$, $\mathcal{D}$ represent images, texts, and distribution data, respectively. Given the mobility feature $\mathbf{m}_{i}$ of a grid and its corresponding observations in other modalities $\mathbf{I}_{i}$, $\mathbf{t}_{i}$, $\mathbf{d}_{i}$ , we encode them into normalized embeddings by modality-specific encoders: $f(\mathbf{m}_{i})\in\mathbb{R}^{d}$ and $g(\mathbf{x}_{i})\in\mathbb{R}^{d}$ where $\mathbf{x}\in\{\mathbf{I},\mathbf{t},\mathbf{d}\}$. Specifically, $\mathit{f}$ is a graph neural network, whereas $\mathit{g}$ are pretrained foundation models for texts and images (see Section 3.2), and an MLP for demographic categorical features. The encoders are optimized with the objective
@@
[…]

### v4

# MoRA: Mobility as the Backbone for Geospatial Representation Learning at Scale

@@ Abstract
Representation learning of geospatial locations remains a core challenge in achieving general geospatial intelligence, with increasingly diverging philosophies and techniques. While Earth observation paradigms excel at depicting locations in their physical states, we propose that a location’s full characterization requires grounding in both its physical attributes and its internal human activity pattern, the latter being particularly crucial for understanding its human-centric functions. We present MoRA, a human-centric geospatial framework that leverages a human mobility graph as its core backbone to fuse various data modalities, aiming to learn embeddings that represent the socio-economic context and functional role of a location. MoRA achieves this through the integration of spatial tokenization, GNNs, and asymmetric contrastive learning to align 100M+ POIs, massive remote sensing imagery, and structured demographic statistics with a billion-edge mobility graph, ensuring the three auxiliary modalities are interpreted through the lens of fundamental human dynamics. To rigorously evaluate the effectiveness of MoRA, we construct a benchmark dataset composed of 9 downstream prediction tasks across social and economic domains. Experiments show that MoRA, with four input modalities and a compact 128-dimensional representation space, achieves superior predictive performances than state-of-the-art models by an average of 12.9%. Echoing LLM scaling laws, we further demonstrate the scaling behavior in geospatial representation learning. We open-source code and pretrained models at: https://github.com/ylzhouchris/MoRA.
@@ 1 Introduction
At the heart of Geospatial Artificial Intelligence (GeoAI) lies a fundamental challenge: the translation of a geographic location—a complex, multifaceted entity defined by physical features, human activity, and latent relationships—into a dense, low-dimensional vector that a machine can process and reason with (Chen et al., 2025; Zhang et al., 2024; Mai et al., 2025). Such a vector, or “representation”, enables models to understand spatial patterns, predict socio-economic outcomes, and track dynamic changes. However, the path toward creating these representations has split into two dominant paradigms, each attempting to answer different fundamental questions about the nature of a place. This divergence shapes every subsequent decision, from data acquisition and model architecture to the very definition of what a location embedding should represent.
The first major paradigm centers on Earth observation. It seeks to describe the “physical” state of locations and build a consistent digital twin of the planet’s surface. Such digital representations provide richer geographic priors than raw coordinates, offering location context that improves downstream tasks such as fine-grained image classification(Mac Aodha et al., 2019; Cole et al., 2023). A standout example within this paradigm is Google’s AlphaEarth (Brown et al., 2025), which ingests petabytes of raw sensor data from a multitude of sources, i.e., optical satellites, radar, LiDAR, and regularizes this heterogeneous information into a unified representation. Other works learns embeddings by aligning coordinates with visual imagery that delineates the physical appearance of locations, either geo-tagged photos or satellite imagery, using contrastive learning (Vivanco et al., 2023; Mai et al., 2023a; Klemmer et al., 2025).
The human‑centric paradigm stands out, in sharp contrast, as the second major stream, built on the premise that a location’s meaning is defined more by the intricate patterns of human activity that unfold within it. Existing studies have employed various types of human‑related data, such as human mobility data (Zhou et al., 2023; Yong and Zhou, 2024) and demographics (Wen and Zhou, 2024), to model socio‑economic dynamics and predict outcomes including housing prices, crime patterns, and consumption levels (Xu and Zhou, 2024; Zhang et al., 2022). However, developing an effective approach, especially grounded by a robust and principled rationale, to utilizing inherently multimodal geospatial data remains a significant and unresolved challenge. There is also a lack of comprehensive benchmarks for human-centric tasks, making it difficult to evaluate the inference capabilities of location representations.
To address these challenges, we claim that the true, latent meaning of a location is primarily defined by its functional relationships with other locations, as revealed through the patterns of human movement, instead of its intrinsic attributes. The theoretical justification is explicitly drawn from an analogy to the functioning of Large Language Models (LLMs), where the semantic meaning of a word (a “token”) is not inherent but emerges from its context—the other words with which it frequently co-occurs in vast text corpora. Similarly, inspired by Vision Transformers (ViT), we discretize continuous geographic space into cells, treating these discrete grid cells as “spatial tokens” and human mobility—sequences of movements from one cell to another—as the ”sentences” that provide context. A cell’s representation is thus profoundly enriched by the latent co-occurrence structures revealed in the sequence-like human movement patterns.
To this end, we propose MoRA, a framework that leverages mobility as the backbone for geospatial representation learning at scale. By using a human mobility graph as the central structural backbone for multimodal data alignment, MoRA ensures that all other data modalities are interpreted through the lens of these fundamental human dynamics. Our contributions can be summarized as follows:
  • We emphasize mobility’s fundamental role as the “syntax of geospatial space”, which gives functional meaning to region tokens by structuring them in sequences, and propose MoRA, a framework that uses a solid relational backbone (mobility) enriched by functional (POIs), social (demographics), and physical (imagery) attributes to learn uniquely comprehensive, humen-centric location representations.
  • Drawing a parallel to the large-scale training paradigms in LLMs, we apply and validate MoRA using a massive mobility graph with extensive long-range links, yielding high-quality, transferable region representations at the national scale. We also provide, to our knowledge, the first empirical evidence of scaling laws in geospatial representation learning.
@@
  • We construct a benchmark evaluation dataset of 9 downstream socio-economic prediction tasks. Extensive experiments show that our model achieves significantly superior performance—improving by 12.9% on average—across all tasks.
  • We open-source our methodology, a benchmark dataset for human-centric representation evaluation, and a distilled version of MoRA presented as a simple, privacy-preserving, ready-to-use utility (coordinates in, vectors out) for national-scale geospatial inference.
@@ 2 Related Works
Representation of the physical states of locations.
Several studies encode the natural and built environmental features of locations into machine-interpretable representations, enabling analysis of natural phenomena and monitoring of environmental changes (Brown et al., 2025; Klemmer et al., 2025). Such approaches often rely on geo-tagged and satellite images, or more broadly on Earth observation data (Yin et al., 2019; 2021). CSP (Mai et al., 2023a) learns embeddings from unlabeled imagery with a contrastive objective, but is pretrained, fine-tuned, and tested on the same datasets (iNat2018 (Van Horn et al., 2018) and fMoW (Christie et al., 2018)), without demonstrating generalizability to other tasks. SatCLIP (Klemmer et al., 2025) and GeoCLIP (Vivanco et al., 2023) both train location encoders via contrastive learning on coordinate-image pairs, differing mainly in encoder architectures and pretraining data. Despite showing success in specific tasks such as worldwide image geo-localization, such visual models may miss crucial non-visual context.
Pretrained foundation model for human-centric geospatial inference.
Geospatial data is inherently diverse and multimodal, spanning maps (Balsebre et al., 2024; Choudhury et al., 2025), texts (Li et al., 2022; 2023; Feng et al., 2024), images (Fan et al., 2023; Huang et al., 2023), and graphs (Wu et al., 2022; Wang and Li, 2017), all widely employed to advance the understanding of urban areas. To leverage this diversity, multimodal region representation learning has emerged as a paradigm that integrates cross-modal information in a self-supervised manner to learn task-agnostic regional embeddings (Yan et al., 2024; Zhang et al., 2021; 2022; Yong and Zhou, 2024). However, such models are often trained separately for each city and show substantial performance drop when transferred across cities—indicative of sensitivity to domain shifts and of gaps in learning city-agnostic geospatial priors. Following the success of pretrained Foundation Models (FMs) in language and vision domains, there is increasing interest in foundation models for geospatial artificial intelligence (GeoAI) (Mai et al., 2024; Zhang et al., 2024). ReFound (Xiao et al., 2024) is an early attempt that distills knowledge from general FMs while adding domain-specific objectives, but its pretraining data, i.e., POIs and satellite image data from five major Chinese cities, are limited in scope and modality diversity compared with ours. Conceptually closest to our work is the Population Dynamics Foundation Model (PDFM) (Agarwal et al., 2024), which encodes postal codes and counties across the United States by leveraging human activity signals such as regional busyness and search trends. Despite advances in scope, it still falls short in modality richness and diversity. More importantly, it fuses modalities through simple concatenation and aggregates information based solely on a static neighborhood graph defined by geographic proximity. This approach misses the dynamic, non-local, and higher-level correlations between regions.
@@
A straightforward way to represent geolocations is to analytically map 2D/3D coordinates into higher-dimensional vectors that preserve geographical distances. Space2Vec (Mai et al., 2020) encodes locations using multi-frequency sinusoid functions in a 2D Euclidean space. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) encodes continuous 3D coordinates via Fourier input mapping. To model Earth’s curvature, Sphere2Vec (Mai et al., 2023b) and Siren (Rußwurm et al., 2024) use position encoders that capture spherical distances between coordinates. These learning-free position encoders are typically followed by a neural network, trained in a supervised manner on task-specific data.
@@ 3 Methodology
The framework learns versatile region representations by combining spatial tokenization, Graph Neural Networks (GNNs), and asymmetric CLIP-based contrastive learning. It adopts a mobility-as-backbone architecture that aligns diverse geospatial modalities to a mobility anchor: Points of Interest (POIs) as texts, satellite imagery as visual data, and demographic distributions as tabular categorical histograms. Figure 1 illustrates the proposed methodological framework.
Figure 1: The methodological framework of MoRA.
@@ 3.1 Mobility as the Backbone
Geographic location provides a precise anchor for aligning multimodal data. Prior region embedding methods (Wu et al., 2024; Klemmer et al., 2025) often fuse modalities via coordinates, focusing solely on imagery, which limits the incorporation of relational signals. Human activity routinely spans geographic boundaries via transport or digital networks, and mobility data—represented as links in a spatial graph—captures these non-local patterns. We propose a mobility graph as the backbone for multimodal fusion and alignment in geospatial representation learning.
@@
Raw spatial interactions are sparse and misaligned in time and space, making coordinate-based graphs too fragmented to reflect major flows. Inspired by Vision Transformers (ViT) (Khan et al., 2022), which partition images into patches, we grid geographic space and build graphs at the cell level. We adopt the H3 grid (Uber Technologies, Inc., ), which minimizes distortion compared to Geohash and Google S2 (Bohui and Weifeng, 2024), offering balanced global coverage. At level 6 ($\sim$36 km² per cell), we construct a nationwide mobility graph from real-world transaction data as presented in the “Graph Construction” section in Figure 1.
@@
We utilize LINE (Large-scale Information Network Embedding) (Tang et al., 2015) to preprocess the full mobility graph and generate 128-dimensional embeddings for each H3 cell. We select LINE embeddings with second-order proximity due to their superior performance in empirical evaluation. Paired with the graph sampling strategy described below, this approach effectively preserves full-graph information with minimal computation.
@@
Graph Neural Networks (GNNs) have proven to be effective for capturing relationships between nodes, taking into account high-order dependencies. GNNs operate through a message-passing mechanism, where each node updates its representation by aggregating information from its neighbors. To encode a graph $\mathcal{G}$ using $L$ layers of GNNs, the $l^{th}$ layer embedding of a node $i$ is computed as follows:
@@
where $\mathcal{N}_{i}$ refers to the neighborhood of the node $i$, and $\textit{AGG}(\cdot)$ is the aggregation function, which can vary across different models. We use LightGCN (He et al., 2020) as the mobility graph encoder. It simplifies the traditional GNN architecture by removing feature transformation and nonlinear activation, and only keeping the most essential component: neighborhood aggregation. In LightGCN, node features are initialized as learnable parameters, and $L$ layers of propagation yield $L+1$ embeddings per node, namely $(e^{(0)},e^{(1)},\dots,e^{(L)})$:
@@
We initialize node features with pre-encoded LINE embeddings and feed a mobility subgraph into the encoder. This subgraph is extracted from the full graph using a top-$k$ sampling strategy based on a percentage threshold, ensuring that the links with the highest flows are retained for each node. This strategy aims to address the ”long-tail” distribution patterns commonly seen in geospatial studies, where a small number of regions experience very high flows, while the majority of regions exhibit significantly lower flows that have minimal impact on the overall graph structure. We adopt a top sampling ratio of 10% for the main experiments.
@@ 3.2 Auxiliary Multimodality Encoder
For geospatial inference, choosing fundamental input modalities that are (1) effective for various human-centric downstream tasks in both social and economic domains, and (2) complementary to each other, is critical. Drawing on spatial data science and urban studies, and supported by prior work, we select three auxiliary modalities to complement the mobility backbone: Points of Interest (POI), demographic statistics, and satellite imagery.
Different unimodal encoders are selected to encode auxiliary modalities. (1) Point-of-Interest (POI): We encode both POI categorical statistics and textual information with the BGE-m3 language model (Chen et al., 2024; Cheng et al., 2025), averaging all POI embeddings within each H3 cell to obtain the grid embedding. (2) Satellite imagery: We use an off-the-shelf pretrained RemoteClip (Liu et al., 2024) encoder to produce grid-level image embeddings. (3) Demographic Statistics: Categorical demographic tabular data is encoded with an MLP-based encoder.
@@ 3.3 Multimodality Alignment with Mobility
We train with CLIP-inspired contrastive learning scheme: all four modalities are projected into a shared space, with mobility as the anchor modality to which the others are aligned. We construct modality tuples $(\mathcal{M},\mathcal{I},\mathcal{T},\mathcal{D})$ for all grids in the study area, where $\mathcal{M}$ indicates mobility, and $\mathcal{I}$, $\mathcal{T}$, $\mathcal{D}$ represent images, texts, and distribution data, respectively. Given the mobility feature $\mathbf{m}_{i}$ of a grid and its corresponding observations in other modalities $\mathbf{I}_{i}$, $\mathbf{t}_{i}$, $\mathbf{d}_{i}$ , we encode them into normalized embeddings by modality-specific encoders: $f(\mathbf{m}_{i})\in\mathbb{R}^{d}$ and $g(\mathbf{x}_{i})\in\mathbb{R}^{d}$ where $\mathbf{x}\in\{\mathbf{I},\mathbf{t},\mathbf{d}\}$. Specifically, $\mathit{f}$ is a graph neural network, whereas $\mathit{g}$ are pretrained foundation models for texts and images (see Section 3.2), and an MLP for demographic categorical features. The encoders are optimized with the objective
@@
which maximizes the similarity between a grid’s mobility features and its corresponding POI, image, or demographic features, while minimizing similarity to those of other grids in the same batch $\left(\mathbf{m}_{i},\mathbf{x}_{i}\right)_{i=1}^{N}$, and a symmetric loss $\mathcal{L}_{\mathcal{X},\mathcal{M}}$ doing the other way around. $\langle\cdot,\cdot\rangle$ represents the dot product operation and $\tau$ is a temperature parameter adjusting the smoothness of the logits in the softmax.